# Kibble-Zurek behavior in a topological phase transition with a quadratic band crossing

### v1

@@ III.2 Disordered system
Then, we switch to another perspective: real-space analysis in disordered systems. Following the previous studies, correlation length can be characterized by the real-space distribution of the local Chern marker (LCM) when the system is weakly disordered, where the LCM $c(\mathbf{r})$ is calculated via Prodan et al. (2010); Bianco and Resta (2011); Ulčakar et al. (2020)
@@
Here $\hat{P}(\mathbf{r})$ is the local projector to the unit cell at $\mathbf{r}$, $\hat{P}_{\mathrm{val}}$ is the projector to the valence band (the lower half spectrum), and $\hat{x},\hat{y}$ are the position operators. The LCM has the properties that it is uniform and equal to the Chern number in clean systems. As long as the disorder is weak and takes the thermodynamic limit, the mean value of LCM is still the Chern number in its clean corresponding, and the local deviation from the mean value shows a domain-like configuration, whose amplitude is proportional to the disorder strength, but its shape is independentUlčakar et al. (2020). Besides, if there are enough disorder samples, the disorder-averaged normalized autocovariance function converges to a stable limit. Thus, the disorder-averaged normalized autocovariance function calculated from a weakly disordered system can be regarded as an intrinsic character of its clean corresponding.
We consider the disorder is added on the time-reversal breaking hopping term, $V\to V_{\mathrm{dis}}(r)$, for $V$ is the only relevant parameter that controls the transition, with an independent uniform distribution $V_{\mathrm{dis}}(r)\in[V-\delta V,V+\delta V]$ with $\delta V=0.05$. The real-space distribution of the LCM in a certain disordered sample with periodic boundary conditions is illustrated in Fig.3(a) for several values of $V$, ranging from topological regions with $C=1$, across the critical point to another topological phase with $C=-1$. We can see that the sizes of domains reach their maximum value (about half the whole system size) at the transition, which indicates the divergence of correlation length in the thermodynamic limit.
We identify the correlation length $\xi_{r}$ in the real space to the distance where the disorder-averaged autocovariance function of the LCM distribution drops to zero App . Fig.3(b) shows that $\xi_{r}$ exhibits power-law scaling as $V$ approaches the critical point in the intermediate correlation length regime, i.e. $64<\xi_{r}<N/$ 6. That is because we are dealing with a finite-size lattice model, the maximal $\xi_{r}$ can not exceed the system size (infrared limit), and the minimal $\xi_{r}$ must saturate to the order of one lattice constant (ultraviolet limit). Only the intermediate regime suffers less from these two limits.
As the system size increases, the estimated scaling exponent approaches $1/2$, which is consistent with the results obtained from the Berry curvature in the clean system. From the above discussion, we find that in a 2D checkerboard model with quadratic band crossing, the critical phenomena of the topological phase transition are present in both momentum space and real space, with a power-law exponent equal to half of the results of linear band crossings.
@@ IV Quench dynamics
In this section, we investigate the dynamical behavior of topological phase transitions in the 2D checkerboard lattice. The system is quenched across the topological phase transition by linearly changing the complex NN hoppings strength $V$ with time $t$, expressed as $V(t)=-1/2+t/\tau$, where $\tau$ denotes the quench time, which corresponds to the change in the Chen number from $1$ to $-1$ as shown by the arrow in the Fig.1(b).
We still work in two perspectives and focus on the freeze-out time $t_{\mathrm{f}}$ and unfrozen correlation length $\xi(t_{\mathrm{f}})$.
IV.1 Clean system
@@ IV.1.1 Freeze-out time $t_{\mathrm{f}}$
Figure 4: (a). The dynamics of non-equilibrium Hall response during the quench. These curves collapse in a rescaled time coordinate. We identify the freeze-out time $t_{\mathrm{f}}$ as the time when the Hall response crosses zero, $\tilde{\Sigma}_{xy}(\hat{t}=t_{\mathrm{f}})=0$. The inset shows curves in a non-rescale $\hat{t}$ coordinate. (b). Freeze-out time with different quench time in a log-log plot, $t_{\mathrm{f}}$ scales with approximately $\sim\tau^{1/2}$. Here, the NNN hopping strength is set to $J^{\prime}=1$.
To explore the temporal scaling in the dynamical process, we choose to study the Hall response during the non-equilibrium process Kuo et al. (2021). The non-equilibrium Hall response is given by the integral of the Berry curvature $F_{\mathrm{b}}(\mathbf{k},t)$ weighted by the occupation probability $P_{\mathrm{b},\mathbf{k}}(t)$ in the upper (lower) band Caio et al. (2016); Hu et al. (2016):
$\begin{split}\tilde{\Sigma}_{xy}(t)=\frac{1}{2\pi}\sum_{\mathrm{b}}\int_{\mathrm{FBZ}}\text{d}^{2}\mathbf{k}P_{\mathrm{b},\mathbf{k}}(t)F_{\mathrm{b}}(\mathbf{k},t)\end{split}$ (5)
Here $\mathrm{b}\in\left\{\mathrm{val},\mathrm{con}\right\}$ refer to the valence and conduction band. We computed the evolution of the Hall response during the quench process, as depicted in Fig.4(a). The Hall response deviates from its original quantized value and relaxes to a new quantized value after the quench. By transforming the coordinates with $\hat{t}=t-t_{\mathrm{c}}$, we set the phase transition time $t_{\mathrm{c}}=\tau/2$ as the zero point and the freeze-out time $t_{\mathrm{f}}$ is defined as the time at which the Hall response relaxes to the average of its initial and final values. We plot the relationship between the freeze-out time $t_{\mathrm{f}}$ and the quench time $\tau$ in the log-log scale and found that it follows a power-law relationship, with the corresponding exponent close to $1/2$. This matches the prediction of the KZ mechanism for linear quenching with $z\nu=1$, $t_{\mathrm{f}}\sim\tau^{\frac{z\nu}{1+z\nu}}\simeq\tau^{1/2}$.
@@ IV.1.2 Unfrozen correlation length $\xi_{k}(t_{\mathrm{f}})$
We subsequently investigate the spatial scalings. The KZ mechanism predicts that at the freeze-out time, the correlation length follows a power-law relationship. By obtaining the evolved wave functions during the quenching process, we calculated the corresponding Berry curvature and extracted the correlation length. Fig.5(a) illustrates the radial distribution of the Berry curvature at corresponding freeze-out time $f(k,{t_{\mathrm{f}}})$ under different quench time. We observed that the width of the peak decreases with increasing quench time. Furthermore, the relationship between the extracted correlation length and the quench time is plotted for $J_{1}^{\prime}=-J_{2}^{\prime}=J^{\prime}=0.5,1,1.5$ and $J=1$. We found power-law relationships between the correlation length and quench time for different correlation lengths, with corresponding critical exponents very close to $1/4$. From the critical exponent of the correlation length at equilibrium, we obtain $\nu\simeq 1/2$. The power-law behavior of the freeze-out time gives the dynamic exponent $z\simeq 2$. According to the KZ mechanism, the unfrozen correlation length scales as $\xi(t_{\mathrm{f}})\sim\tau^{\nu/(1+z\nu)}\simeq\tau^{1/4}$. Therefore, our results are consistent with the KZ prediction. This critical exponent with a quadratic band crossing is half of the linear band crossing Ulčakar et al. (2020); Sun et al. (2022).
Figure 5: (a). Radial distribution of the Berry curvature at the freeze-out time $t_{\mathrm{f}}$ with $J^{\prime}=1$. (b). Correlation length $\xi_{k}$ at the freeze-out time obtained by the Berry curvature has a scaling with quench time $\tau$, $\xi\sim\tau^{1/4}$.
IV.2 Disordered system
@@ IV.2.1 Freeze-out time $t_{\mathrm{f}}$
It is not straightforward to generalize the Hall response to the disordered systems. We need real-space analysis to show analogs to the formation of topological defects. Therefore, we switch back to explore the LCM distribution.
The distribution of the LCM in quench dynamics can be calculated by replacing the projector onto the valence band $\hat{P}_{\mathrm{val}}$ with the projector onto the instant occupied subspace. The evolution of the LCM profile during the quench process with $\tau=10$ and $\tau=100$ under weak disorder is depicted in Fig.6(a). At the beginning of the slow quench, their LCM profiles are similar since they both follow the adiabatic evolution. For longer quench time, the system enters into the freeze-out regime later and holds a larger domain size after quench.
Figure 6: (a). The spatial distribution of the local Chern marker during a fast ($\tau=10$) and a slow ($\tau=100$) quench with size $N=50$. The slower the quench, the larger the post-quench domain size. (b). The dynamics of correlation length in real space in a rescaled coordinate. Curves collapse near the transition point. The freeze-out time $t_{\mathrm{f}}$ can be determined by the time delay of the trend of divergence in $\xi_{r}$ while crossing the transition. The inset shows the same data in a non-rescaled coordinate. Each data is an average from $10$ disorder samples. (c). Left panel: freeze-out time has a power-law scaling with the quench time, and the power-law exponent is close to $1/2$ for different system sizes. Right panel: estimated exponents with different system sizes. (d). Left panel: correlation length at the freeze-out time shows a scaling behavior, and the power-law exponent is close to $1/4$ for different system sizes. Right panel: estimated exponents with different system sizes. Here, the NNN hopping strength is set to $J^{\prime}=1$ and the disorder strength to $\delta V=0.05$.
Like before, we extract the evolution of the correlation length $\xi_{r}$ over different quenching time, as shown in Fig.6(b). During the evolution, the correlation length $\xi_{r}$ adiabatically follows the instant ground-state value initially. Near the transition, the system enters into the freeze-out regime. Although the correlation length still increases, the trend of divergence occurs later than the critical point, so it shows a maximum after the transition. Here, we claim this time delay could be identified as the freeze-out time $t_{\mathrm{f}}$ (We verify that this method is also applicable to clean systems as discussed in Appendix App ). After that, the correlation length is stabilized around a constant larger than the ground-state value.
@@ IV.2.2 Unfrozen correlation length $\xi_{r}(t_{\mathrm{f}})$
Similarly, we extract the unfrozen correlation length $\xi_{r}(t_{\mathrm{f}})$ App from the LCM at the freeze-out time $t_{\mathrm{f}}$ to analyze. We plot the relationship between unfrozen correlation length $\xi_{r}(t_{\mathrm{f}})$ and the quench time on a log-log scale in the Fig.6(d). We find that $\xi_{r}(t_{\mathrm{f}})$ scales with $\tau$ in an exponent of approximately $1/4$. Correspondingly, the critical exponent of the correlation length at equilibrium yields $\nu\simeq 1/2$, and the power-law behavior of the freeze-out time gives the dynamic exponent $z\simeq 2$. According to KZ mechanism, the unfrozen correlation length is expected to scale as $\xi(t_{\mathrm{f}})\sim\tau^{\nu/(1+z\nu)}\simeq\tau^{1/4}$. Thus, our findings are in agreement with the KZ prediction as observed in the clean system.
Therefore, in the checkerboard lattice with quadratic band crossing, when quenching across the topological phase transition, we observe temporal and spatial power-law behavior analogous to the Kibble-Zurek (KZ) mechanism in both momentum space and real space. The temporal power-law exponent is the same as that of linear band crossing, while the spatial power-law exponent is half of that.
V Outlook on higher-order band crossing
The critical exponent obtained from the quadratic band crossing differs from that of the previously studied linear band crossing. Therefore, it remains uncertain whether the KZ power-law scaling behaviors persist in systems with higher-order dispersion relations. If so, the corresponding critical exponent and its relationship with the band dispersion need further investigation. Accordingly, we can write down a general low-energy two-band effective model with higher-order crossing in momentum space, and explore the KZ behavior via methods mentioned in clean systems.
@@ V Outlook on higher-order band crossing
$\begin{split}H_{p}(k,\phi)&=\eta_{p}k^{p}\cos(p\phi)\sigma_{x}+\eta_{p}k^{p}\sin(p\phi)\sigma_{y}+M\sigma_{z},\end{split}$ (6)
where $k$ and $\phi$ represent the distance and polar angle relative to band-crossing momentum, respectively and integer $p$ denotes its crossing order, with $p=1$ representing linear band crossing and $p=2$ representing quadratic band crossing, and so forth. The topological phase transition occurs at $M=0$ with the change in Chern number before and after the transition given by $C(M\textgreater 0)-C(M\textless 0)=p$. Based on this Hamiltonian, we can first compute the Berry curvature in the ground state and obtain the correlation length. We numerically verify that the correlation length in ground state satisfies $\xi\sim|M-M_{\mathrm{c}}|^{-\nu}$, where $\nu$ is approximately $1/p$.
Similarly to the previous analysis, we quench across the topological phase transition, measure the Hall response, and extract the power-law relationship between the freeze-out time and the quench time. We find that temporal scaling is independent of the order of band crossing with the corresponding exponent close to $1/2$. Similarly, by measuring the time evolution of the Berry curvature, we obtain the correlation length at the freeze-out time under different quench time, which also satisfies a power-law relationship. We observe that the corresponding exponents are distinct and close to $(2p)^{-1}$.
Figure 7: (a). Correlation length in the ground state diverges in a power $\nu\simeq 1/3$ for $p=3$ and $\nu\simeq 1/4$ for $p=4$ as the system approaches the phase transition. (b). Freeze-out time has a power-law scaling with the quench time, $t_{\mathrm{f}}\sim\tau^{1/2}$ with $p=3$ and $p=4$. (c). Correlation length at the freeze-out time shows a scaling behavior for $p=3$, $\xi(t_{\mathrm{f}})\sim\tau^{1/6}$ and $p=4$, $\xi(t_{\mathrm{f}})\sim\tau^{1/8}$.

### v2

@@ III.2 Perspective II: Disordered system
Then, we switch to the other perspective: real-space analysis in disordered systems. Following the previous studies, correlation length can be characterized by the real-space distribution of the local Chern marker (LCM) when the system is weakly disordered, where the LCM $c(\mathbf{r})$ is calculated via Prodan et al. (2010); Bianco and Resta (2011); Ulčakar et al. (2020)
@@
Here $\hat{P}(\mathbf{r})$ is the local projector to the unit cell at $\mathbf{r}$, $\hat{P}_{\mathrm{val}}$ is the projector to the valence band (the lower half spectrum), and $\hat{x},\hat{y}$ are the position operators.
If the system is clean, the LCM is uniform and equal to the Chern number in the ground state. While in a disordered system, as long as the disorder is weak, the mean value of LCM of the ground state is still the Chern number of its original clean counterpart in the thermodynamic limit Oliveira and Chen (2024). In a specific disorder sample, the local deviation from the mean value shows a domain-like configuration (see Fig.3(a)), whose amplitude is proportional to the disorder strength, but its shape is independent Ulčakar et al. (2020). Besides, if there are enough disorder samples, the disorder-averaged normalized autocorrelation function of LCM converges to a certain limit (see Appendix for more details). Therefore, the disorder-averaged normalized autocorrelation function calculated from a weakly disordered system can be regarded as an intrinsic character of its clean counterpart, and the correlation length determined by this method is a intrinsic quantity of the topological system.
We consider the weak disorder is added on the time-reversal-breaking hopping term, $V\to V_{\mathrm{dis}}(\mathbf{r})$, with an independent uniform random distribution $V_{\mathrm{dis}}(\mathbf{r})\in[V-\delta V,V+\delta V]$ with $\delta V=0.05$. The real-space distribution of the LCM in a specific disordered sample is illustrated in Fig.3(a) for several values of $V$, ranging from topological regions with $C=1$, across the critical point to another topological phase with $C=-1$. At the transition, the sizes of domains reach their maxima for about the whole system size, which indicates the divergence of correlation length in the thermodynamic limit.
We identify the correlation length in the real space $\xi_{r}$ to the distance where the disorder-averaged autocorrelation function of the LCM distribution drops to zero Ulčakar et al. (2020); Sun et al. (2022); Molignini et al. (2023). Fig.3(b) shows that $\xi_{r}$ exhibits power-law scaling as $V$ approaches the critical point in the intermediate correlation length regime, i.e. $64<\xi_{r}<N/$ 6. That is because we are dealing with a finite-size lattice model, the maximal $\xi_{r}$ can not exceed the system size (infrared limit), and the minimal $\xi_{r}$ must saturate to the order of one lattice constant (ultraviolet limit). Only the intermediate regime suffers less from these two limits.
As the system size increases, the estimated scaling exponent $\nu$ approaches $1/2$, which is consistent with the results obtained from the Berry curvature in the clean system. From the discussion above, we find that in a 2D checkerboard model with a quadratic band crossing, the critical phenomena of the topological phase transition are present in both momentum space and real space, with a power-law exponent equal to half of the results of linear band crossings.
@@ IV Quench dynamics
In this section, we investigate the KZ behavior of topological phase transitions in the 2D checkerboard lattice. The system is quenched across the topological transition by linearly changing the complex NN hoppings strength $V$ with time $t$, expressed as $V(t)=-1/2+t/\tau$, where $\tau$ denotes the quench time, which corresponds to the change in the Chern number from $1\left(t=0\right)$ to $-1\left(t=\tau\right)$ as shown by the arrow in the Fig.1(b).
We still work in two perspectives and focus on the freeze-out time $t_{\mathrm{f}}$ and the unfrozen correlation length $\xi(t_{\mathrm{f}})$.
IV.1 Perspective I: Clean system
@@ IV.1.1 Freeze-out time $t_{\mathrm{f}}$
Figure 4: (a). The dynamics of non-equilibrium Hall response during the quench with $J^{\prime}=1$. These curves collapse in a rescaled time coordinate. We identify the freeze-out time $t_{\mathrm{f}}$ as the time when the Hall response crosses zero, $\tilde{\Sigma}_{xy}(\hat{t}=t_{\mathrm{f}})=0$. The inset shows curves in a non-rescale $\hat{t}$ coordinate. (b). Freeze-out time with different quench time in a log-log plot, $t_{\mathrm{f}}$ scales with approximately $\sim\tau^{1/2}$.
To explore the temporal scaling in the dynamical process, we choose to study the Hall response during the non-equilibrium process Kuo et al. (2021). The non-equilibrium Hall response is given by the integral of the Berry curvature $F_{\mathrm{b}}(\mathbf{k},t)$ weighted by the occupation probability $P_{\mathrm{b}}(\mathbf{k},t)$ in all instantaneous bands Caio et al. (2016); Hu et al. (2016):
$\begin{split}\tilde{\Sigma}_{xy}(t)=\frac{1}{2\pi}\sum_{\mathrm{b}}\int_{\mathrm{FBZ}}\text{d}^{2}\mathbf{k}~{}P_{\mathrm{b}}(\mathbf{k},t)F_{\mathrm{b}}(\mathbf{k},t).\end{split}$ (5)
Here $\mathrm{b}\in\left\{\mathrm{val},\mathrm{con}\right\}$ refers to the instantaneous valence and conduction band. $P_{\mathrm{b}}(\mathbf{k},t)$ equals to the overlap between the eigenstates of the instantaneous Hamiltonian and the time-evolved wave function during the quench, while $F_{\mathrm{b}}(\mathbf{k},t)$ is calculated using the corresponding instantaneous eigenstates via Eq.3. We computed the evolution of the Hall response during the quench process, as depicted in Fig.4(a). The Hall response deviates from its original quantized value and relaxes to a new quantized value after the quench. By transforming the coordinates with $\hat{t}=t-t_{\mathrm{c}}$, we set the phase transition time $t_{\mathrm{c}}=\tau/2$ as the zero point and the freeze-out time $t_{\mathrm{f}}$ is defined as the time at which the Hall response relaxes to the average of its initial and final values. We plot the relationship between the freeze-out time $t_{\mathrm{f}}$ and the quench time $\tau$ in the log-log scale and found that it follows a power-law relationship, with the corresponding exponent close to $1/2$. This matches the prediction of the KZ mechanism for linear quenching with $z\nu\simeq 1$, $t_{\mathrm{f}}\sim\tau^{\frac{z\nu}{1+z\nu}}\simeq\tau^{1/2}$.
@@ IV.1.2 Unfrozen correlation length $\xi_{k}(t_{\mathrm{f}})$
We subsequently investigate the spatial scalings. The KZ mechanism predicts that at the freeze-out time $t_{\mathrm{f}}$, the correlation length $\xi_{k}(t_{\mathrm{f}})$ follows a power-law scaling with $\tau$. By obtaining the evolving wave functions during the quenching process, we calculated the corresponding Berry curvature and extracted the correlation length. Fig.5(a) illustrates the radial distribution of the Berry curvature at corresponding freeze-out time $f(k,{t_{\mathrm{f}}})$ under different quench time. We observed that the width of the peak decreases with increasing quench time. Furthermore, the relationship between the extracted correlation length and the quench time is plotted for $J^{\prime}=0.5,1,1.5$. We found power-law relationships between the correlation length and quench time for different correlation lengths, with corresponding critical exponents very close to $1/4$. From the critical exponent of the correlation length at equilibrium, we obtain $\nu\simeq 1/2$. The power-law behavior of the freeze-out time gives the dynamic exponent $z\simeq 2$. According to the KZ mechanism, the unfrozen correlation length scales as $\xi(t_{\mathrm{f}})\sim\tau^{\nu/(1+z\nu)}\simeq\tau^{1/4}$. Therefore, our results are consistent with the KZ prediction. This critical exponent with a quadratic band crossing is half of the linear band crossing Ulčakar et al. (2020); Sun et al. (2022).
Figure 5: (a). Radial distribution of the Berry curvature at the freeze-out time $t_{\mathrm{f}}$ with $J^{\prime}=1$. (b). Correlation length $\xi_{k}$ at the freeze-out time obtained by the Berry curvature has a scaling with quench time $\tau$, $\xi_{k}(t_{\mathrm{f}})\sim\tau^{1/4}$.
IV.2 Perspective II: Disordered system
@@ IV.2.1 Freeze-out time $t_{\mathrm{f}}$
When it comes to disordered systems, we need real-space analysis to show analogs to the formation of topological defects. Therefore, we switch back to explore the LCM distribution.
The distribution of the LCM in quench dynamics can be calculated by replacing the projector onto the valence band $\hat{P}_{\mathrm{val}}$ with the projector onto the instantaneous occupied subspace. The evolution of the LCM profile during the quench process with $\tau=10$ and $\tau=100$ under weak disorder is depicted in Fig.6(a). At the beginning of the slow quench, their LCM profiles are similar since they both follow the adiabatic evolution. For longer quench time, the system enters into the freeze-out regime later and holds a larger domain size after quench.
Figure 6: (a). The spatial distribution of the LCM during a fast ($\tau=10$) and a slow ($\tau=100$) quench with size $N=50$. The slower the quench, the larger the post-quench domain size. (b). The dynamics of correlation length in real space in a double-rescaled coordinate. Curves collapse near the transition point. The freeze-out time $t_{\mathrm{f}}$ can be determined by the time delay of the trend of divergence in $\xi_{r}$ while crossing the transition. The inset shows the same data in a non-rescaled coordinate. Each data is an average from $10$ disorder samples. (c). Left panel: freeze-out time has a power-law scaling with the quench time, and the power-law exponent is close to $1/2$ for different system sizes. Right panel: estimated exponents with different system sizes. (d). Left panel: correlation length at the freeze-out time shows a scaling behavior, and the power-law exponent is close to $1/4$ for different system sizes. Right panel: estimated exponents with different system sizes. Here, the NNN hopping strength is set to $J^{\prime}=1$ and the disorder strength to $\delta V=0.05$.
Like before, we extract the evolution of the correlation length $\xi_{r}$ over different quenching time, as shown in Fig.6(b). During the evolution, the correlation length $\xi_{r}$ adiabatically follows the instantaneous ground-state value initially. Near the transition, the system enters into the freeze-out regime. Although the correlation length still increases, the trend of divergence occurs later than the critical point, so it shows a maximum after the transition. Here, we claim this time delay could be identified as the freeze-out time $t_{\mathrm{f}}$ (we verify that this method is also applicable to momentum-space analysis in clean systems as discussed in Appendix). After that, the correlation length is stabilized around a constant larger than the ground-state value.
@@ IV.2.2 Unfrozen correlation length $\xi_{r}(t_{\mathrm{f}})$
Similarly, we extract the unfrozen correlation length $\xi_{r}(t_{\mathrm{f}})$ from the LCM at the freeze-out time $t_{\mathrm{f}}$ to analyze. We plot the relationship between unfrozen correlation length $\xi_{r}(t_{\mathrm{f}})$ and the quench time on a log-log scale in the Fig.6(d). We find that $\xi_{r}(t_{\mathrm{f}})$ scales with $\tau$ in an exponent of approximately $1/4$. Correspondingly, the critical exponent of the correlation length at equilibrium yields $\nu\simeq 1/2$, and the power-law behavior of the freeze-out time gives the dynamic exponent $z\simeq 2$. According to KZ mechanism, the unfrozen correlation length is expected to scale as $\xi(t_{\mathrm{f}})\sim\tau^{\nu/(1+z\nu)}\simeq\tau^{1/4}$. Thus, our findings are in agreement with the KZ prediction as observed in the clean system.
Therefore, in the checkerboard lattice with a quadratic band crossing, when quenching across the topological phase transition, we observe temporal and spatial power-law behavior analogous to the Kibble-Zurek (KZ) mechanism in both momentum space and real space. The temporal power-law exponent is the same as that of the linear band crossing, while the spatial power-law exponent is half of that.
V Outlook on higher-order band crossings
The critical exponent obtained from the quadratic band crossing differs from that of the previously studied linear band crossings. Therefore, it remains uncertain whether the KZ power-law scaling behaviors persist in systems with higher-order band crossings. If so, the corresponding critical exponent and its relationship with the band dispersion need further investigation. Accordingly, we can write down a general low-energy two-band effective model with higher-order crossing in momentum space, and explore the KZ behavior via methods mentioned in clean systems.
@@ V Outlook on higher-order band crossings
$\begin{split}\hat{H}_{p}(k,\phi)&=\eta_{p}k^{p}\cos(p\phi)\hat{\sigma}_{x}+\eta_{p}k^{p}\sin(p\phi)\hat{\sigma}_{y}+M\hat{\sigma}_{z},\end{split}$ (6)
where $k$ and $\phi$ represent the distance and polar angle relative to band-crossing momentum, respectively and integer $p$ denotes its crossing order, with $p=1$ representing the linear band crossing and $p=2$ representing the quadratic band crossing, and so forth. The topological phase transition occurs at $M=0$ with the change in Chern number between before and after the transition given by $C(M>0)-C(M<0)=p$. We should clarify that this effective description is only valid around the band-closing momentum near the transition. As we are studying the slow quench physics close to the transition, the distinction between this and the original lattice model is not important here.
Based on this Hamiltonian, we can first compute the Berry curvature in the ground state and obtain the correlation length. We numerically verify that the correlation length in ground state satisfies $\xi_{k}\sim|M-M_{\mathrm{c}}|^{-\nu}$, where $\nu$ is approximately $1/p$.
Similarly to the previous analysis, we quench across the topological phase transition, measure the Hall response, and extract the power-law relationship between the freeze-out time and the quench time. We find that temporal scaling is independent of the order of the band crossing with the corresponding exponent close to $1/2$. Similarly, by measuring the time evolution of the Berry curvature, we obtain the correlation length at the freeze-out time under different quench time, which also satisfies a power-law relationship. We observe that the corresponding exponents are distinct and close to $(2p)^{-1}$.
Figure 7: (a). Correlation length in the ground state diverges in a power $\nu\simeq 1/3$ for $p=3$ and $\nu\simeq 1/4$ for $p=4$ as the system approaches the phase transition. (b). Freeze-out time has a power-law scaling with the quench time, $t_{\mathrm{f}}\sim\tau^{1/2}$ with $p=3$ and $p=4$. (c). Correlation length at the freeze-out time shows a scaling behavior for $p=3$, $\xi_{k}(t_{\mathrm{f}})\sim\tau^{1/6}$ and $p=4$, $\xi_{k}(t_{\mathrm{f}})\sim\tau^{1/8}$.